OpenStudy (anonymous):

Use the equation given to find f "(pi/3) f(x) = sec(x) f"(pi/3)= ?

OpenStudy (anonymous):

f(x)=sec(x) f'(x)=sec(x)tan(x) f"(x)=sec(x)[tan²(x)+sec²(x)] with x=(pi/3) you can calculate what you need